OpenStudy (anonymous):

(2x10^-3)x(2x10^8) Multiply & write the answer w/ exponents

OpenStudy (anonymous):

for this question keep in mind that if you have x^a *x^b = x^a+b alright you could only add if the bases are the same so we got (2^1+1)*10^(-3+8) = (2^2)*(10^5) :)